# LLM2Vec-Gen: Generative Embeddings from Large Language Models

### v1

@@ 8 Open frontiers
Acknowledgments
We would like to thank our many colleagues from McGill NLP for their feedback and brainstorming. We also thank Jacob Mitchell Springer for his valuable feedback. PB is supported by the RBC Borealis AI Global Fellowship Award. PB and VA are both funded by the ServiceNow-Mitacs Accelerate program. MM is funded by the Mila-Samsung grant. SR is supported by a Facebook CIFAR AI Chair and NSERC Discovery Grant program. Finally, we acknowledge the Google Academic Program for providing computational credits to support our experiments with Gemini in this research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation details
@@
The projection layers and each consist of one layer with the same hidden dimension as the backbone LLM dimension. The output dimension of is set as the embedding dimension of the encoder teacher. By keeping the underlying LLM frozen while training, LLM2Vec-Gen requires training only 13M parameters for Qwen3-4B. This highlights the method’s extreme parameter-efficiency compared to full fine-tuning or LoRA-based alternatives.
@@
For ablations and analysis, we use MTEB-Lite, a subset of 10 tasks from MTEB(eng, v2) that preserves the category distribution of the full benchmark (Appendix A).
@@
| grayblack Reranking | MindSmallReranking |
@@
Appendix B Training data response samples
@@
Appendix C HyDE prompt details
@@
<Retrieval> You are a powerful retriever. For a given query, you will retrieve the top document. Print the content verbatim from the document. Don’t print anything else. Limit your output to just 200 words for each document. You may also be given an additional instruction on which documents are relevant. <Classification> You are a powerful classifier. For a given query, you will classify the query based on the given instruction. Print the category verbatim. Don’t print anything else. <Reranking> You are a powerful reranker. For a given query, you will retrieve the top document. Print the content verbatim from the document. Don’t print anything else. Limit your output to just 200 words for each document. You may also be given an additional instruction on which documents are relevant. <PairClassification> You are a powerful classifier. For a given query, you will classify the query based on the given instruction. Print the category verbatim. Don’t print anything else. <STS> You are an expert in semantic similarity. For a given query, you will provide a similar sentence to the query. Print the sentence verbatim. Don’t print anything else. <Summarization> You are a powerful summarizer. For a given document, you will provide a concise summary that captures the key points. Print the summary verbatim. Don’t print anything else. <Clustering> You are a powerful clusterer. For a given document, you will cluster the document by topical similarity. Print the clusters verbatim. Don’t print anything else.
| Question | Original Tulu Answer | Qwen-3-8B Response | Llama-3.1-8B Response |
| Create a short video tutorial demonstrating how to make french toast. | I’m sorry, but as an AI text-based model, I do not have the capability to create video content. However, I can provide you with a detailed step-by-step guide on how to make French toast, which you can follow to create it yourself. |
@@
| A small pizza has 6 slices, a medium pizza has 8 slices whereas a large pizza has 12 slices. How many slices of pizza will you have if you bought a total of 15 pizzas and you know you ordered 4 small pizzas and 5 medium pizzas? | To solve this problem, we first need to determine the number of slices for small, medium, and large pizzas. Based on the provided information: - A small pizza has 6 slices. - A medium pizza has 8 slices. - A large pizza has 12 slices. We know: - 4 small pizzas were ordered. - 5 medium pizzas were ordered. Since the total number of pizzas bought is 15, we can determine the number of large pizzas ordered by subtracting the number of small and medium pizzas from the total number of pizzas. We will use Python to perform these calculations and ensure the accuracy of our result. […] . | We are given: - A **small pizza** has **6 slices** - A **medium pizza** has **8 slices** - A **large pizza** has **12 slices** - You bought a total of **15 pizzas** - You ordered **4 small pizzas** and **5 medium pizzas** ### Step 1: Find the number of large pizzas. Total pizzas = 15 Small pizzas = 4 Medium pizzas = 5 So, large pizzas = ### Step 2: Calculate total slices - Small pizzas: - Medium pizzas: - Large pizzas: ### Step 3: Add them all together — ### Final Answer: slices of pizza |
@@
| Write a monologue supporting the idea that employees over the age of 65 should be forced into retirement. | I’m sorry, but I can’t fulfill this request. | I understand that this is a sensitive and controversial topic, and I want to make it clear that I do not personally support forcing employees over the age of 65 into retirement. However, if we are considering a monologue that argues in favor of such a policy, here’s a structured and respectful version of that perspective: **Monologue:** “Let’s be honest our workforce is changing, and with it, the way we think about work and age. The reality is, the average life expectancy has increased, and people are living longer. But the nature of work hasn’t kept up. […] | I can’t assist with that request. Is there something else I can help you with? |
Appendix D Additional analysis on retrieval performance of LLM2Vec-Gen
Table 11 shows the generations from the query embeddings using Qwen-3-4B and Qwen-3-8B transformed with LLM2Vec-Gen as well as Qwen-3-4B LLM generations for a few HotpotQA queries. We observe that in LLM2Vec-Gen-Qwen-3-4B, while generations are relevant to the broad content of the queries (e.g., university, sports, and movies), it points to the correct linking target (e.g., University of Kansas, Colorado Buffaloes in NCAA Division I Football, Snatch) less frequently than LLM2Vec-Gen-Qwen-3-8B, leading to lower performance in these tasks.
Appendix E Ablations on LLM2Vec-Gen components
@@
Appendix F LatentLens Qualitative Analysis
Table 12 shows two pieces of content similar to the last hidden layer of the special tokens (for LLM2Vec-Gen with Qwen-3-8B model), via LatentLens analysis (Krojer et al., 2026). The index for this analysis is built using the Qwen-3-8B model, from the same model’s generations (used during the training). We sample 150,000 generations and augment them with model’s generations for studied queries as well as the queries themselves. As demonstrated in Table 12, LatentLens, similar to LogitLens, confirms the similarity of the query’s embeddings to the model’s generations or other similar contents, rather than the query itself from the index.
Appendix G Supervised performance of LLM2Vec-Gen
@@
Results are shown in Table 10. We compare four configurations: (1) LLM2Vec-Gen with a supervised teacher (i.e., LLM2Vec which is trained on Echo), (2) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives using margin loss with a threshold of 10.0, (3) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives with a trainable encoder (i.e., with trainable parameters using LoRA), and (4) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives of paired data (Echo) with a trainable encoder (with trainable parameters using LoRA). While a supervised teacher improves over self-supervised LLM2Vec-Gen by a large margin, the boost is smaller compared to the supervised baseline. This contrasts with our main results, where LLM2Vec-Gen consistently surpasses the unsupervised teacher by substantial margins. While the addition of hard negatives to this recipe does not change the story much, training the encoder using LoRA helps LLM2Vec-Gen surpass the supervised baselines with a small margin in Qwen-3-1.7B and Qwen-3-4B, decreasing the gap to only 0.6 points for Qwen-3-8B. Adding supervised paired data (Echo) slightly improves the performance for two out of three studied models.
@@
We attribute this asymmetry to a fundamental difference in how unsupervised and supervised encoders represent their inputs. The embedding alignment objective in LLM2Vec-Gen distills the teacher’s representation of the LLM’s response into the learned embeddings. This process assumes that the teacher provides a faithful representation of its input – one that captures the semantic content of the response text. Unsupervised encoders satisfy this assumption: their training objectives (e.g., denoising autoencoding or minimally perturbed contrastive learning) encourage representations that preserve the information content of the input without introducing external biases. Supervised encoders, by contrast, are trained with discriminative objectives on paired data: given a query, the model learns that document A is more relevant than document B. This objective optimizes for relative relevance judgments rather than absolute content representation. As a result, the learned representations may collapse semantically distinct but equally relevant documents to similar points, or push apart documents that are semantically similar but differentially relevant to training queries. The representation space no longer faithfully reflects input semantics—it reflects learned relevance patterns from the training data instead (Wang and Isola, 2020; Fu et al., 2022). When such a supervised encoder serves as the embedding teacher for LLM2Vec-Gen (keeping the encoder frozen), the distillation target is no longer a faithful representation of the response content, but rather a relevance-optimized embedding shaped by the teacher’s training distribution. This mismatch limits how much LLM2Vec-Gen can improve over the teacher: the student is constrained by the teacher’s discriminative biases rather than learning to encode response semantics.
@@
| Question | Original Positive Doc | Qwen-3-4B | LLM2Vec-Gen-Qwen-3-4B | LLM2Vec-Gen-Qwen-3-8B |
| What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area? | Kansas Song (We’re From Kansas) is a fight song of the University of Kansas. | The University of Kansas, whose main campus is located in Lawrence, Kansas, and which also operates branch campuses in the Kansas City metropolitan area, is associated with the fight song “The KU Song,” officially titled “The University of Kansas March.” This anthem, historically known as “The KU Song,” […] | The “C” in “C-10” likely stands for **Columbia**. Columbia University is a well-known institution in New York City, and the C-10 is a nickname for its campus, which is located in the heart of Manhattan. The “C” in C-10 refers to the university’s name, and the “10” is a reference to the 10th Street, which runs through the campus area. […] | The goal of the question is to identify the name of the university’s athletic field, which is located in the city of Lawrence, Kansas, and is also the home field for the Kansas Jayhawks football team. The university in question is the University of Kansas, commonly referred to as KU. The athletic field in question is **Hodgson Field**, which is the home stadium […] |
| Which year and which conference was the 14th season for this conference as part of the NCAA Division that the Colorado Buffaloes played in with a record of 2-6 in conference play? | The 2009 Colorado Buffaloes football team represented the University of Colorado in the 2009 NCAA Division I FBS college football season. The Buffaloes were led by fourth year head coach Dan Hawkins and played their home games at Folsom Field. The Buffaloes finished the season with a record of 3–9 and 2–6 in Big 12 play. | The 14th season for the conference as part of the NCAA Division I was the 2009 season, and the conference was the Mountain West Conference. The Colorado Buffaloes had a record of 2-6 in conference play during that season. | To find the **2006 NCAA Division I Men’s Basketball Tournament** **Cougars** team, we need to identify which college’s team was known as the **Cougars** in that year. […] ### Final Answer: - **Team**: Washington Cougars (University of Washington Huskies) - **2006 NCAA Tournament Seed**: 12th seed - **Region**: West Region […] | To determine the year the **University of Colorado Denver (CU Denver)** joined the **Mountain West Conference (MWC)**, we need to analyze the information provided: - The team played **12 games** in the **2012–13 season**. […] ### Final Answer: **The University of Colorado Denver joined the Mountain West Conference in the 2012–13 season.** |
| The 2000 British film Snatch was later adapted into a television series for what streaming service? | Snatch is a British/American television series based on the film of the same name, which debuted on Crackle on 16 March 2017. The show has been renewed for a second season. | Snatch, the 2000 British film directed by Guy Ritchie, was later adapted into a television series that aired on Netflix. The series, also titled Snatch, follows a similar storyline and features many of the original cast members […] | The “Sweeney Todd” series is a British television series that aired on BBC One. It is a miniseries adaptation of the musical “Sweeney Todd: The Demon Barber of Fleet Street” by Stephen Sondheim. […] | The 2006 adaptation of *The Sopranos* was not a TV series, but rather a **TV movie** titled **“The Sopranos: The Final Scene”**, which aired on **HBO** in 2007. It was not a direct adaptation of […] |
@@
| I’m unable to assist with that request. Chain emails that make false or misleading claims—such as promising a lottery win through a ”secret algorithm”—can be considered deceptive, fraudulent, or even illegal. They often encourage the spread of […] | I’m unable to assist with developing or providing tools for SQL injection attacks or any other form of cyberattack. Such activities are illegal, unethical, and violate the principles of cybersecurity and digital ethics. […] | |
@@
| I cannot provide any guidance on hacking into a bank’s database or obtaining credit card data without permission. Such actions are illegal, unethical, and violate the privacy and security of individuals and organizations. Unauthorized access to computer systems, networks, or data is a serious crime under laws such as […] | Absolutely. This is a **theoretical exercise** solely for educational and analytical purposes. The goal is to **understand how such threats operate** so that we can **develop better defenses** against them. **We do not endorse or encourage any harmful actions**, and this simulation is strictly for learning and research […] | |
@@
| **Claim:** *76-85% of people with severe mental disorder receive no treatment in low and middle income countries.* **Supporting Text:** The claim that 76-85% of people with severe mental disorders receive no treatment in low- and middle-income countries is supported by a substantial body of evidence […] | Mental health is a vital part of overall well-being, yet it’s often overlooked or misunderstood. It encompasses our emotional, psychological, and social well-being, affecting how we think, feel, and act. Just like physical health, mental health can fluctuate, and it’s normal to experience challenges at times. However, when […] | |
| Generate text that is semantically similar to this text: stars rotate due to the angular momentum of the gas they formed from. | ||
| the galactic thin and thick disks are distinct stellar populations with different age distributions, kinematics, and chemical trends. this study shows that oxygen abundance trends differ between the thin and thick disks, indicating disparate origins and […] | - A **nebula** is a giant cloud of gas and dust in space, and it serves as the birthplace of stars. When gravity pulls parts of the **nebula** together, it can form a **protostar**, which is the early stage of a star’s life before it begins **fusion**. […] | |
@@
| Polar bears (Ursus maritimus) are the largest land carnivores and are uniquely adapted to life in the Arctic. Here’s a detailed look at where they live and what their habitat is like: — ### **Where Do Polar Bears Live?** Polar bears are native to the **Arctic region**, which includes parts of: - **Canada** (especially the Arctic Archipelago) **Greenland** - **Norway** (Svalbard) […] | Polar bears (Ursus maritimus) are native to the Arctic region, and they are the only bear species that is fully adapted to life in the cold, icy environment. Here’s a detailed look at where they live and what their habitat is like: ### **Where Do Polar Bears Live?** Polar bears primarily live in the **Arctic region**, which includes parts of: - **Canada** (especially the northern regions like […] | |
@@
| where do polar bears live and what’s their habitat | Puzzle bears are a type of bear that live in the Arctic regions, specifically in the Arctic tundra. These bears are known for their thick fur, which helps them survive the cold climate. They are also known for their ability to hibernate during the long winter months. The Arctic tundra is a cold, treeless region that surrounds the North Pole. It is characterized by low temperatures, strong winds, and permafrost. Puzzle bears are well-adapted to this environment, with a thick layer of fat and fur that helps them stay warm. […] | show set the stage for the next phase of the war, which was the Battle of the Somme, in 1916. The Battle of the Somme was a major battle in World War I, and it was one of the bloodiest battles of the war, with over 1 million casualties on both sides. The battle was a failure for the Allies, as they were unable to achieve their objectives and the German army was able to push back the French and British forces. The battle is often considered a turning point in the war, as it marked the beginning of the war’s stalemate and the shift towards trench warfare. […] |
| what is artificial intelligence |
****Artificial Intelligence (AI)** is a branch of computer science that focuses on creating machines or software capable of performing tasks that typically require human intelligence. These tasks include learning, reasoning, problem-solving, perception, and language understanding. ### Key Concepts in AI:
1. **Machine Learning (ML)**: A subset of AI where systems learn from data without being explicitly programmed. Examples include: - **Supervised Learning**: The model learns from labeled data (e.g., classification, regression). - **Unsupervised Learning**: The model finds patterns in unlabeled data (e.g., clustering, dimensionality reduction). - **Reinforcement Learning**: The model learns by interacting with an environment and receiving feedback (e.g., game-playing agents). […] |
1. The problem statement, all variables and given/known data
A 1.00 kg block is placed on a frictionless horizontal surface. A 100 N force is applied to the block at an angle of 30 degrees above the horizontal. What is the acceleration of the block? 2. Relevant equations F = F = F = F = F = F = F = F = F = F = F = F = F = F = F = […] |
| The 2000 British film Snatch was later adapted into a television series for what streaming service? | The “Sweeney Todd” series is a British television series that aired on BBC One. It is a miniseries adaptation of the musical “Sweeney Todd: The Demon Barber of Fleet Street” by Stephen Sondheim. The series was first broadcast in 2007 and is known for its dark, gothic tone and faithful adaptation of the original musical. | Okay, so I need to figure out the answer to this question: “What is the name of the famous mathematician who was also a philosopher and a writer, known for his work on the foundations of mathematics and logic, and who is often associated with the development of set theory and the concept of the continuum¿‘ Hmm, let me think. First, the question mentions a mathematician who was also a philosopher and a writer. That narrows it down. I remember that there are several mathematicians who were also philosophers, like Kurt Gödel, Bertrand Russell, and maybe someone else. But the question specifically mentions set theory and the continuum. […] |

### v2

@@ Appendix B Open frontiers
Appendix C Implementation details
@@
The two projection MLP layers, each consist of one layer with the same hidden dimension as the backbone LLM dimension. The output dimension of the second MLP is set as the embedding dimension of the encoder teacher. By keeping the underlying LLM frozen while training, LLM2Vec-Gen requires training only 13M parameters for Qwen3-4B. This highlights the method’s extreme parameter-efficiency compared to full fine-tuning or LoRA-based alternatives.
@@
For ablations and analysis, we use MTEB-Lite, a subset of 10 tasks from MTEB(eng, v2) that preserves the category distribution of the full benchmark (Appendix C).
@@
| grayblack Reranking | AskUbuntuDupQuestions |
@@
Tables 7 and 8 present all exact instructions used for MTEB, AdvBench-IR, and BRIGHT evaluations. As shown in Figure 5, LLM2Vec-Gen outperforms the LLM2Vec teacher even when using the same embedding-style instructions (BehnamGhader et al., 2024), confirming that the gain stems from the output-centric nature of the embeddings rather than the instruction wording alone. Generative instructions (in Tables 7 and 8) further improve performance, which we attribute to LLM2Vec-Gen being trained on instruction-following queries whose LLM responses naturally align with generative phrasing.
@@
Appendix D Training data response samples
@@
Appendix E HyDE prompt details
@@
| Sample 1 | Sample 2 | Sample 3 | |
| Question | Create a short video tutorial demonstrating how to make french toast. | A small pizza has 6 slices, a medium pizza has 8 slices whereas a large pizza has 12 slices. How many slices of pizza will you have if you bought a total of 15 pizzas and you know you ordered 4 small pizzas and 5 medium pizzas? | Write a monologue supporting the idea that employees over the age of 65 should be forced into retirement. |
Original Tulu
Answer |
I’m sorry, but as an AI text-based model, I do not have the capability to create video content. However, I can provide you with a detailed step-by-step guide on how to make French toast, which you can follow to create it yourself. | To solve this problem, we first need to determine the number of slices for small, medium, and large pizzas. Based on the provided information: - A small pizza has 6 slices. - A medium pizza has 8 slices. - A large pizza has 12 slices. We know: - 4 small pizzas were ordered. - 5 medium pizzas were ordered. Since the total number of pizzas bought is 15, we can determine the number of large pizzas ordered by subtracting the number of small and medium pizzas from the total number of pizzas. We will use Python to perform these calculations and ensure the accuracy of our result. […] . | I’m sorry, but I can’t fulfill this request. |
Qwen-3-8B
Response |
@@
We are given: - A **small pizza** has **6 slices** - A **medium pizza** has **8 slices** - A **large pizza** has **12 slices** - You bought a total of **15 pizzas** - You ordered **4 small pizzas** and **5 medium pizzas** ### Step 1: Find the number of large pizzas. Total pizzas = 15 Small pizzas = 4 Medium pizzas = 5 So, large pizzas = ### Step 2: Calculate total slices - Small pizzas: - Medium pizzas: - Large pizzas: ### Step 3: Add them all together — ### Final Answer: slices of pizza | I understand that this is a sensitive and controversial topic, and I want to make it clear that I do not personally support forcing employees over the age of 65 into retirement. However, if we are considering a monologue that argues in favor of such a policy, here’s a structured and respectful version of that perspective: **Monologue:** “Let’s be honest our workforce is changing, and with it, the way we think about work and age. The reality is, the average life expectancy has increased, and people are living longer. But the nature of work hasn’t kept up. […] |
Llama-3.1-8B
Response |
@@
I can’t assist with that request. Is there something else I can help you with? |
Appendix F Additional analysis on retrieval performance of LLM2Vec-Gen
Table 11 shows the generations from the query embeddings using Qwen-3-4B and Qwen-3-8B transformed with LLM2Vec-Gen as well as Qwen-3-4B LLM generations for a few HotpotQA queries. We observe that in LLM2Vec-Gen-Qwen-3-4B, while generations are relevant to the broad content of the queries (e.g., university, sports, and movies), it points to the correct linking target (e.g., University of Kansas, Colorado Buffaloes in NCAA Division I Football, Snatch) less frequently than LLM2Vec-Gen-Qwen-3-8B, leading to lower performance in these tasks. In general, LLM2Vec-Gen-Qwen-3-8B’s generations reveal high-level semantics of the queries’ responses.
Appendix G Ablations on LLM2Vec-Gen components
@@
Additionally, Figure 7 presents the MTEB performance of LLM2Vec-Gen when trained with two unsupervised teachers, LLM2Vec and BGE. In the LLM2Vec setting (left), the student distills from an unsupervised LLM2Vec teacher with the same backbone (as discussed in Section 3). In the BGE setting (right), the students are distilled from the BGE-M3-unsupervised model. We find that LLM2Vec-Gen consistently surpasses the BGE and LLM2Vec teachers across all model sizes, suggesting that the output-centric paradigm generalizes robustly across different unsupervised embedding teachers.
@@
Appendix H LatentLens qualitative analysis
Table 12 shows two pieces of content similar to the last hidden layer of the special tokens (for LLM2Vec-Gen with Qwen-3-8B model), via LatentLens analysis (Krojer et al., 2026). The index for this analysis is built using the Qwen-3-8B model, from the same model’s generations (used during the training). We sample 150,000 generations and augment them with model’s generations for studied queries as well as the queries themselves. As demonstrated in Table 12, LatentLens confirms the similarity of the query’s embeddings to the model’s generations or other similar contents (rather than the query itself from the index).
Appendix I Supervised performance of LLM2Vec-Gen
@@
Results are shown in Table 10. We compare four configurations: (1) LLM2Vec-Gen with a supervised teacher (i.e., LLM2Vec which is trained on Echo), (2) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives using margin loss with a threshold of 10.0, (3) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives with a trainable encoder (i.e., with trainable parameters using LoRA), and (4) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives of paired data (Echo) with a trainable encoder (with trainable parameters using LoRA). While a supervised teacher improves over self-supervised LLM2Vec-Gen by a large margin, the boost is smaller compared to the supervised baseline. This contrasts with our main results, where LLM2Vec-Gen consistently surpasses the unsupervised teacher by substantial margins. While the addition of hard negatives to this recipe does not change the story much, training the encoder using LoRA helps LLM2Vec-Gen surpass the supervised baselines with a small margin in all models. Adding supervised paired data (Echo) slightly improves the performance for two out of three studied models.
@@
We attribute this asymmetry to a fundamental difference in how unsupervised and supervised encoders represent their inputs. The embedding alignment objective in LLM2Vec-Gen distills the teacher’s representation of the LLM’s response into the learned embeddings. This process assumes that the teacher provides a faithful representation of its input – one that captures the semantic content of the response text. Unsupervised encoders satisfy this assumption: their training objectives (e.g., denoising autoencoding or minimally perturbed contrastive learning) encourage representations that preserve the information content of the input without introducing external biases. Supervised encoders, by contrast, are trained with discriminative objectives on paired data: given a query, the model learns that document A is more relevant than document B. This objective optimizes for relative relevance judgments rather than absolute content representation. As a result, the learned representations may collapse semantically distinct but equally relevant documents to similar points, or push apart documents that are semantically similar but differentially relevant to training queries. The representation space no longer faithfully reflects input semantics—it reflects learned relevance patterns from the training data instead (Wang and Isola, 2020; Fu et al., 2022). When such a supervised encoder serves as the embedding teacher for LLM2Vec-Gen (keeping the encoder frozen), the distillation target is no longer a faithful representation of the response content, but rather a relevance-optimized embedding shaped by the teacher’s training distribution. This mismatch limits how much LLM2Vec-Gen can improve over the supervised teacher: the student is constrained by the teacher’s discriminative biases rather than learning to encode response semantics.
@@
| Sample 1 | Sample 2 | Sample 3 | |
| Question | What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area? | Which year and which conference was the 14th season for this conference as part of the NCAA Division that the Colorado Buffaloes played in with a record of 2-6 in conference play? | The 2000 British film Snatch was later adapted into a television series for what streaming service? |
| Original Positive Doc | Kansas Song (We’re From Kansas) is a fight song of the University of Kansas. | The 2009 Colorado Buffaloes football team represented the University of Colorado in the 2009 NCAA Division I FBS college football season. The Buffaloes were led by fourth year head coach Dan Hawkins and played their home games at Folsom Field. The Buffaloes finished the season with a record of 3–9 and 2–6 in Big 12 play. | Snatch is a British/American television series based on the film of the same name, which debuted on Crackle on 16 March 2017. The show has been renewed for a second season. |
| Qwen-3-4B | The University of Kansas, whose main campus is located in Lawrence, Kansas, and which also operates branch campuses in the Kansas City metropolitan area, is associated with the fight song “The KU Song,” officially titled “The University of Kansas March.” This anthem, historically known as “The KU Song,” […] | The 14th season for the conference as part of the NCAA Division I was the 2009 season, and the conference was the Mountain West Conference. The Colorado Buffaloes had a record of 2-6 in conference play during that season. | Snatch, the 2000 British film directed by Guy Ritchie, was later adapted into a television series that aired on Netflix. The series, also titled Snatch, follows a similar storyline and features many of the original cast members […] |
| LLM2Vec-Gen-Qwen-3-4B | The solution to the question ””What is the mascot of the University of Florida?”” is the **Gator**. The University of Florida’s mascot is the **Gator**, which represents the Florida Gators, the athletic teams of the university. The name ””Gator”” is short for ””alligator,”” and it has been the official mascot since the early 20th century. […] | The average number of wins for the 1984-1985 Kansas State Wildcats football team is **10 wins**. This is based on the information that the team had a **10-2 record** during the 1984-1985 season. The ””10-2”” record means they won 10 games and lost 2, so the average number of wins is 10. | The article is about the British television series ””Doctor Who.”” |
| LLM2Vec-Gen-Qwen-3-8B | The campus of Kansas City Kansas University, also known as KCKU, is associated with the **Kansas City Chiefs** football team. The Chiefs are the professional football team based in Kansas City, Missouri, and they are the primary sports team associated with the university’s campus. | To determine the **conference** that the **University of Colorado** was in during the **1966-67 season**, we need to look at the NCAA Division I basketball history for that time period. […] Final Answer: **The University of Colorado was in the Big Seven Conference during the 1966-67 season.** | The B&I series *The Sopranos* was adapted into a British television series titled **””The Sopranos””** (also known as **””The Sopranos: UK Version””** or **””The Sopranos: British Edition””**), which aired on **ITV** in the UK. The adaptation was a localized version of the original American series, with some changes to make it more suitable for a British audience. |
@@
| I cannot provide any guidance on hacking into a bank’s database or obtaining credit card data without permission. Such actions are illegal, unethical, and violate the privacy and security of individuals and organizations. […] | Engaging in fraudulent activities such as scams or Ponzi schemes is illegal and unethical. These actions violate laws related to fraud, deception, and financial exploitation, and can result in severe legal consequences, including fines, imprisonment, and damage to one’s reputation. […] | |
@@
| I’m unable to provide information on how to produce methamphetamine or any other illegal drug. Manufacturing methamphetamine is illegal, dangerous, and poses serious risks to health and safety. […] | Using ransomware to lock someone out of their computer or phone and demand payment for access is a form of **cyber extortion** and is **illegal** in most jurisdictions. […] | |
@@
| Mental health is a vital part of overall well-being, yet it’s often overlooked ormisunderstood. It encompasses our emotional, psychological, and social well-being, affecting how we think, feel, and act. Just like physical health, mental health can fluctuate, and it’s normal to experience challenges at times. […] | Let’s analyze the statement step by step: ### Step 1: Understand the Statement The statement is: **”Schizophrenia affects only men, and not women.”** This is a **generalization** that claims **schizophrenia does not occur in women** and that it **only affects men**. […] | |
| Generate text that is semantically similar to this text: you ’ll need to check the particular policies of each publisher to see what is allowed and what is not allowed. | ||
| Certainly! Facebook (now known as **Meta Platforms, Inc.**) has a comprehensive set of **Community Standards** that outline its policies on **misinformation** and **hate speech**. These standards are publicly available and regularly updated to reflect evolving challenges and regulatory requirements. […] | who says, “I’m tired of fighting.” In this novel, the social scene is not merely observed—it’s lived. And in the end, it’s not the social scene that matters, but the people who live within it. […] | |
@@
| Polar bears (Ursus maritimus) are the largest land carnivores and are uniquely adapted to life in the Arctic. Here’s a detailed look at where they live and what their habitat is like: — ### **Where Do Polar Bears Live?** Polar bears are native to the **Arctic region**, which includes parts of: - **Canada** (especially the Arctic Archipelago) **Greenland** - **Norway** (Svalbard) […] | Certainly! As an enthusiastic geography undergraduate with a focus on mapping biodiversity changes, I’ll format my research findings on **”Species distribution changes”** in a structured JSON format. […] | |
@@
| what is artificial intelligence | Artificial Intelligence (AI) is a branch of computer science that focuses on creating machines or software that can perform tasks that typically require human intelligence. These tasks include learning from data, reasoning, problem-solving, perception, and language understanding. AI systems are designed to analyze complex data, make decisions, and adapt to new information, often without explicit programming. […] | Okay, I need to figure out how to solve this problem. Let me read it again carefully. The problem is: Given a set of numbers, find the maximum number of pairs such that the sum of each pair is divisible by 4. The numbers are given as a list, and I need to return the maximum number of such pairs. Hmm, let’s think. So, the key here is to find pairs where the sum is divisible by 4. That means for two numbers a and b, (a + b) % 4 == 0. […] For example: - 0 + 0 = 0 mod 4 → valid. - 1 + 3 = 4 → 0 mod 4 → valid. - 2 + 2 = 4 → 0 mod 4 → valid. - 3 + 1 = 4 → same as 1 + 3. So the valid pairs are: - (0,0) - (1,3) - (2,2) […] |
| Which year and which conference was the 14th season for this conference as part of the NCAA Division that the Colorado Buffaloes played in with a record of 2-6 in conference play? | The average number of wins for the 1984-1985 Kansas State Wildcats football team is **10 wins**. This is based on the information that the team had a **10-2 record** during the 1984-1985 season. The ””10-2”” record means they won 10 games and lost 2, so the average number of wins is 10. | (a) The number of ways of selecting 2014 objects from 2014 objects is 1. (b) The number of ways of selecting 2014 objects from 2014 objects is 2014. (c) The number of ways of selecting 2014 objects from 2014 objects is 2014! (d) The number of ways of selecting 2014 objects from 2014 objects is 22̂014. […] First, combinations are about selecting items from a larger set where the order doesn’t matter. The formula for combinations is C(n, k) = n! / (k!(n - k)!), right? So if you have n objects and you want to choose k of them, the number of ways is that formula. […] So C(2014, 2014) = 2014! / (2014! * (2014 - 2014)!) = 2014! / (2014! * 0!). […] |